OpenStudy (anonymous):

What is the value of integral (2 to 5): (d(2x-5)^4 / dx) * dx ?

OpenStudy (anonymous):

Do you mean this? \[\int\limits_{2}^{5}\frac{d(2x-5)^4}{dx}dx\]

OpenStudy (cwrw238):

this is integrated using the extended form of the power rule Intregral of ax^m = a x^(m+1) / (m+1)

OpenStudy (anonymous):

Yes @tom982

OpenStudy (cwrw238):

here you are replacing x by 2x-5 so you multiply by 1/2 ( 1 divided by the coefficient of x)

OpenStudy (cwrw238):

oh - you are integrating the derivative of (2x - 5)^4 I misread the question

OpenStudy (anonymous):

Doesn't the integral of the derivative just cancels out leave us with: \[\left[(2x-5)^4 \right]_2^5=624\]

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

Thanks for the medal :)